# Rapid Bayesian identification of sparse nonlinear dynamics from scarce and noisy data

### v1

@@ 2.3 Comparing derivative approximation using weak formulation and finite difference
One of the major challenges in applying SINDy-type algorithms to learn governing equations from data is approximating temporal (for learning ODEs) and spatial (for learning PDEs) derivatives using discrete data in the presence of noise. Conventional finite differences can greatly amplify noise, rendering the dynamics unrecoverable for the STLS algorithm. To overcome this problem, one can employ the weak formulation [13, 14]. This formulation uses Galerkin projection and a compact and smooth test function $\phi(t)$ to “transfer” the derivatives on the data to the test function. For example, the Galerkin projection allows the time derivative on the data $\hat{x}^{(k)}(t)$ to be expressed as
@@
provided that $\phi(t)$ and its derivatives up to $\phi^{(k)}(t)$ have compact support in $t\in[a,b]$. As a result, any noise present in the data is not magnified, leading to a substantial enhancement in the performance of the subsequent regression.
@@
Note that both $L_{\partial t}$ and $L_{I}$ are rectangular matrices with dimension $\tilde{N}\times N$, where $\tilde{N}$ is the number of measurements in the original dataset, whereas $N\leq\tilde{N}$ is the number of data after computing the derivatives. $\tilde{N}-N=2n$ number of data points are usually filtered away near the two ends of the data series due to a lack of data beyond the measured point to compute the corresponding derivative. Also, notice that the scaling of $|L_{\partial t}|:|L_{I}|$ is of order $\mathcal{O}(1/\Delta t)$ in finite difference, whereas that of weak formulation is of order $\mathcal{O}(1)$. This is why, when applied to (5), finite difference amplifies noise, while weak formulation does not.
@@
Therefore, the decision between weak formulation and finite difference involves a compromise between attenuating noise and maintaining signal. The finite difference method does not attenuate any signal from a spectrum standpoint, but it amplifies noise by the magnitude of the sampling rate. Hence, it is more suitable for data sampled at a low sampling rate. Conversely, the weak formulation is a proficient method for attenuating noise and mitigating noise amplification, but it may also suppress signals at high frequencies. It is suitable for data that is sampled at a rate significantly higher than the Nyquist frequency of the signal. In practice, the sampling rate is often limited by the capabilities of the measuring hardware, and problems that push the limit of the equipment are often sampled at a frequency just above the Nyquist frequency of the relevant dynamics. As a rule of thumb, weak formulation usually requires at least 6-12 data points in the integral. Hence, practically, it can only resolve dynamics 6-12 times slower than the sampling frequency. By improving model selection performance in the high noise regime, the Bayesian reformulation presented in this work improves the overall chances of recovering the correct model even when the sampling rate is close to the Nyquist frequency of the dynamic, i.e. in the low-data regime.
@@ 3 Examples
This section will showcase the noise robustness and data efficiency of our Bayesian-SINDy algorithm in recovering the desired equations precisely in comparison with the original SINDy algorithm [1] that is based on Sequentially Thresholded Least Squares (STLS). We will also compare our Bayesian-SINDy algorithm with SparseBayes [24] to show the improvements our Bayesian-SINDy has brought to the Bayesian approaches. Four examples will be used to compare the three algorithms, in which we will consider a polynomial library $\Theta(\mathbf{X})$ up to third order. To compare the three algorithms and to ensure the result is not skewed by individual runs, we will vary the sampling frequency and noise variance for the first three synthetic examples, and iterate the learning process 10000 times to compare the percentage of runs that each algorithm has successfully recovered exactly all the terms in original equation without extra spurious term, i.e. success rate of recovering the original equations. In this work, the use of success rate as the metric is justified as we are focused on the most parsimonious and generalisable equations, which are the original equations, precisely. Furthermore, to avoid miscounting models with the same sparsity pattern but substantially different coefficients, the success rate is only counted when the model coefficient error (MCE = $||\mathbf{w}-\hat{\mathbf{w}}||_{2}/||\hat{\mathbf{w}}||_{2}$, where $\hat{\mathbf{w}}$ is the truth) is below $25\%$. As for the fourth example, we will learn the dynamical equations from real-world measurements. Although a ground-truth model is absent, historical uses of the dataset [34, 35, 31] have shown that we should be expecting the Lotka-Volterra equations.
We should note that in this section, we are not aiming to claim the Bayesian-SINDy to be the most effective method in recovering the correct dynamical equations from data, nor is SINDy (STLS) the most effective methodology among many SINDy extensions. As we shall discuss later in §4.2, we expect more computationally-intensive sampling-based methods to be more robust. However, by comparing SINDy (STLS) with an equivalently simple and computationally efficient method derived from the Bayesian framework, we will demonstrate that a Bayesian approach is more robust in noisy and low data scenario, and thereby encourage a Bayesian perspective to other extensions of SINDy. The comparison with SparseBayes, meanwhile, demonstrates the improvement we have brought to our learning algorithm. To compare the algorithms properly, we have also taken steps to ensure they are performing the best they can with the appropriate hyperparameters. Details of the hyperparameters used for each algorithm and in each example are listed in the appendix.
@@ 3.2 Cubic oscillator
Figure 5: The phase diagram of the cubic nonlinear oscillator system (22, black line) and the system estimated by SINDy (STLS) (blue), Bayesian-SINDy (red) and SparseBayes/RVM (yellow) given the noisy data (crosses) of the system. The noise level are $(a)$ $\sigma_{x}=0.005$ and $(b)$ $\sigma_{x}=0.01$ and the data are sampled at $\Delta t=0.05$ for $t\in[0,5]$. The right panel shows the model equations from each algorithm when $\sigma_{x}=0.01$. In $(a)$, the prediction by SINDy and B-SINDy is so close to truth their differences are practically indiscernible.
Figure 6: The percentage of runs that successfully identifies the cubic nonlinear oscillator equations from data using SINDy (STLS, blue), Bayesian-SINDy (red) and SparseBayes (yellow) after 10000 iterations against the number of data points $\tilde{N}=5/\Delta t$. The noise levels are $(a)$ $\sigma=0.005$ and $(b)$ $\sigma=0.01$.
@@
Figure 7: Cubic oscillator: breakdown of contributions to the total variance of noise $\boldsymbol{\sigma}$ in the dynamics of $(a)$ $x_{1}$ and $(b)$ $x_{2}$ by each term in (9). Thick lines show the variance computed from the statistics of 1000 iterations, and lighter lines (shaded area) show the bound of the theoretical calculation using (9) from each iteration.
@@ 3.3 Lorenz system
Figure 8: The percentage of runs that successfully identifies the Lorenz system equations from data sampled between $t\in[0,2.5]$ using SINDy (STLS, blue), Bayesian-SINDy (red) and SparseBayes (yellow) after 10000 iterations against the number of data points $\tilde{N}=2.5/\Delta t$. The noise levels are $(a)$ $\sigma=0.025$, $(b)$ $\sigma=0.05$, $(c)$ $\sigma=0.1$ and $(d)$ $\sigma=0.2$. The dashed line shows the success rate of recovering the model without the $x_{2}$ term in the equation for $\dot{x}_{2}$. The inset in $(d)$ shows the success rate when the time interval is extended to $t\in[0,10]$.
However, in figure 8$(d)$, in which the noise is $\sigma=0.2$, none of the three algorithms can recover the model exactly. Both SINDy (blue) and SparseBayes (yellow) can only recover the full model sporadically and showed little consistency in the model they produced. In contrast, Bayesian-SINDy consistently recovers a model with only the $x_{2}$ term in equation for $\dot{x}_{2}$ missing (as shown by the dashed red line). The missing $x_{2}$ term in $\dot{x}_{2}$ is particularly difficult to learn because it is highly collinear with the existing $x_{1}$ term in the model. To distinguish the two would require more data in regions in which the two differ significantly, but, as shown in figure 9, the given signal is highly obscured by noise and does not sample those regions sufficiently in the short time interval $t\in[0,2.5]$. In other words, there is simply not enough information in the small time window to infer the $x_{2}$ term in the equation for $\dot{x}_{2}$. From the evidence-maximisation perspective, this difficulty is reflected in figure 1, where $\mathcal{J}(\alpha_{m})>0$ only for a small range of $\alpha_{m}$. In practice, when the data is too low or too noisy, the information content in the data may not be enough to infer the full dynamics. In such cases, one inevitably has to increase the amount of data available by considering a long time interval, say, $t\in[0,10]$. As demonstrated in the inset of figure 8$(d)$, Bayesian-SINDy can eventually recover the full dynamics most of the time, while SINDy remains less robust in this high noise regime. This lack of information, as depicted here in figure 9, also hints at the possibility of further improving data efficiency if we are allowed to sample data in the most informative state. We will discuss this information-theoretical approach further in §4.3.
Figure 9: Learning the dynamics of $\dot{x}_{2}-rx_{1}+x_{1}x_{2}$ in the Lorenz system: the crosses show the values of $\dot{x}_{2}-rx_{1}+x_{1}x_{3}$ from the data, which should match $x_{2}$ (red line). However, $x_{1}$ (yellow line) is highly collinear with $x_{2}$, and it is not obvious which term should fit the data better. Both $x_{1}$ and $x_{2}$ are scaled to best match the data $\dot{x}_{2}-rx_{1}+x_{1}x_{3}$. Here, $\Delta t=0.025$ ($\tilde{N}=100$) and $\sigma=0.2$.
@@ 3.4 Lotka-Volterra equation from Lynx-Hare population data
In this example, we apply the three algorithms to model the population dynamics of Lynx and Hare in Canada, using pelt trade data collected by the Hudson Bay Company between 1900 and 1920 [34, 35]. Assuming the pelts of these two species traded by the Company are proportional to the true population, it may be anticipated that these two populations will approximately adhere to the Lotka-Volterra model for predator-prey dynamics. In fact, this dataset has been used as empirical evidence to support the Lotka-Volterra equations, and has become increasingly popular as a real-life benchmarking dataset among statisticians and data scientists [36, 12, 31]. The primary challenge posed by this dataset is its limited number of data points - specifically, 21 annual data points for each population, resulting in a total of 42 data points. Due to the low sampling rate, the finite difference method is more appropriate, but it will, as one expects, also amplify the noise in the derivative, which we will then mitigate by relying on the noise robustness of the Bayesian framework. The weak formulation was also tried, but the low-pass filtering effect filtered vital signals in the data, leading to subpar learning performance across all methods.
@@
To address the difficulty, we ran a hyperparameter sweep on $\sigma_{x}$, assuming $\alpha=10^{-2}$, and chose the value of $\sigma_{x}\approx 2.7$ by evidence maximisation. Although this technique is similar to the evidence maximisation process in SparseBayes, a notable distinction is that our procedure does not assume a uniform $\boldsymbol{\sigma}$ for all data points. Instead, we calculate it using equation (9) and algorithm 1. The parameters estimated are displayed in table 1, where the Lotka-Volterra model is successfully recovered by Bayesian-SINDy, while the other two struggle to learn a consistent model. Uncertainties in the estimated parameters, also shown in table 1, are also easily recovered from the the posterior distribution. Figure 10 also shows how the uncertainty in the estimated parameter propagates through the dynamics, assuming the same initial condition. Notably, our method was able to estimate the parameter uncertainties without expensive sampling, unlike previous techniques [31].
@@ 4.1 Scaling of likelihood and Occam factor with $N$ and $\boldsymbol{\beta}$
$\mathcal{J}(\mathbf{t}|\mathbf{D},\mathbf{w},\boldsymbol{\beta})=-\frac{N}{2}\ln{2\pi}-\frac{1}{2}\boldsymbol{\beta}\cdot\boldsymbol{\beta}-\frac{1}{2}\boldsymbol{\beta}\cdot(\mathbf{t}-\mathbf{D}\mathbf{w})^{2}.$ (24)
@@
$\mathcal{J}(\mathbf{t}|\mathbf{D},\boldsymbol{\beta},\boldsymbol{\alpha})=\mathcal{J}(\mathbf{t}|\mathbf{D},\mathbf{w},\boldsymbol{\beta})|_{\mathbf{w}=\boldsymbol{\mu}}+\mathcal{J}(\mathbf{w}|\boldsymbol{\alpha})|_{\mathbf{w}=\boldsymbol{\mu}}+\frac{1}{2}\ln{(\det(\boldsymbol{\Sigma}))},$ (25a)
@@
the log-likelihood $\mathcal{J}(\mathbf{t}|\mathbf{D},\mathbf{w},\boldsymbol{\beta})$ and the log-prior $\mathcal{J}(\mathbf{t}|\mathbf{D},\mathbf{w},\boldsymbol{\beta})$ at the maximum a posteriori (MAP) point. The MAP point refers to the parameter values $\mathbf{w}=\boldsymbol{\mu}$ that maximise the posterior. It should be noted that in equation (24), the log-likelihood scales linearly with the magnitude of the inverse noise variance $\boldsymbol{\beta}$ and the number of data points $N$. On the other hand, the remaining components in the log-evidence, also known as the Occam factor, serve as a penalty for models that have an excessive number of parameters. The Occam factor scales with the number of terms $M$ in the model and only has a partial and weak $\ln{N}$ scaling with $N$.
@@
Figure 11: Difference in log-evidence, log-likelihood and the Occam factor (i.e. log-evidence minus log-likelihood) compared to the highest-ranked model of all possible models for $\dot{x}_{3}$ in the Lorenz system. Models are ranked by $(a,c,e)$ evidence and $(b,d,f)$ likelihood. Insets show a zoomed plot of the first few hundred models, and the barcode plots below show the non-zero terms in each model, marked as green (correct) or orange (incorrect). The correct model is always ranked highest by evidence in $(a,e,i)$, and indicated with a star in $(b,d,f)$. Meanwhile, the model selected by highest likelihood contain all the terms, and the models selected by SINDy are shown in the middle. Here, $(a-b)$ $\sigma_{x}=0.2$ and $\tilde{N}=200$, $(c-d)$ $\sigma_{x}=1$ and $\tilde{N}=200$, and $(e-f)$ $\sigma_{x}=1$ and $\tilde{N}=50$.
@@
Figure 11 demonstrates how the scaling of the log-likelihood and the Occam factor plays out in the model selection process. The left column of figure 11 shows the contribution of the log-likelihood (blue) and the Occam factor’s (red) to the log-evidence across all 1024 possible models for $\dot{x}_{3}$ in the Lorenz system when a smaller, second-order polynomial library containing ten terms is considered. For each model, the barcode plot below show the library terms that are non-zero, with the non-zero terms $x_{3}$ and $x_{1}x_{2}$ in the true model are coloured green. The models are sorted by log-evidence, demonstrating a Bayesian perspective on model selection, where the model with maximum evidence is selected as the most parsimonious model for the data. For the same dataset, the right column of figure 11 sort the model based on log-likelihood calculated from the maximum likelihood estimate (MLE), i.e. $\mathbf{w}=\mathbf{\mu}_{\text{MLE}}$, which can be obtained by ordinary least squares. (Note that when $N$ and $\beta$ is large, the MLE and MAP estimates of $\mathbf{\mu}$ converge.) In contrast to the Bayesian perspective on the left, the right column reflects a frequentist perspective where the model with the best fit is selected as the best model for the data. As one would expect, the MLE-based model selection always select the model with all ten terms. To induce sparsity, the frequentist approach STLS introduces a empirical thresholding process. The idea is to remove small coefficient terms that doesn’t contribute much to the likelihood, thereby achieving a compromise between maximising likelihood and sparisty.
@@
When the data is abundant and the noise is minimal (i.e. $N||\boldsymbol{\beta}||\gg M$), models that include the correct terms have significantly higher likelihood and are easily distinguishable from the other models when sorted by likelihood. This is demonstrated by the continuous green band in the barcode plot in 11$(b)$. Phenomenologically, this signifies that all models containing the appropriate terms exhibit comparable levels of likelihood. STLS leverages this clear separation in likelihood and only eliminate irrelevant terms that do not affect the likelihood much in its thresholding process, thereby effectively constraining its search for the most sparse model within this region of large likelihood.
@@
However, as the amount of data is lowered (i.e. smaller $N$), or when the noise increases (i.e. smaller $||\boldsymbol{\beta}||$), the distinct separation of the high likelihood region from the other possible models becomes less evident. In these scenarios, models lacking the relevant terms are more likely to fit the data incidentally, as reflected by the discontinuous green band in barcode plot of figure 11$(d,f)$. STLS may, thus, converge on an incorrect model. Alternatively, STLS may not converge to the most sparse model since there is a non-negligible chance for irrelevant terms also to have coefficients larger than the threshold when $N$ or $\beta$ decreases.

### v2

@@ 2.3 Comparing derivative approximation using weak formulation and finite difference
One of the major challenges in applying SINDy-type algorithms to learn governing equations from data is approximating temporal (for learning ODEs) and spatial (for learning PDEs) derivatives using discrete data in the presence of noise. Conventional finite differences can greatly amplify noise, rendering the dynamics unrecoverable by the STLS algorithm. To overcome this problem, one can employ the weak formulation [13, 14]. This formulation uses Galerkin projection and a compact and smooth test function $\phi(t)$ to “transfer” the derivatives on the data to the test function. For example, the Galerkin projection allows the time derivative on the data $\hat{x}^{(k)}(t)$ to be expressed as
@@
provided that $\phi(t)$ and its derivatives up to $\phi^{(k)}(t)$ are smooth and have compact support in $t\in[a,b]$. As a result, any noise present in the data is not magnified, leading to a substantial enhancement in the performance of the subsequent regression.
@@
Note that both $L_{\partial t}$ and $L_{I}$ are rectangular matrices with dimension ${N}\times\tilde{N}$, where $\tilde{N}$ is the number of measurements in the original dataset, whereas $N\leq\tilde{N}$ is the number of data after computing the derivatives. $\tilde{N}-N=2n$ number of data points are usually filtered away near the two ends of the data series due to a lack of data beyond the measured point to compute the corresponding derivative. Also, notice that the scaling of $|L_{\partial t}|:|L_{I}|$ is of order $\mathcal{O}(1/\Delta t)$ in finite difference, whereas that of weak formulation is of order $\mathcal{O}(1)$. This is why, when applied to (5), finite difference amplifies noise, while weak formulation does not.
@@
Therefore, the decision between weak formulation and finite difference involves a compromise between attenuating noise and maintaining signal. The finite difference method does not attenuate any signal from a spectrum standpoint, but it amplifies noise by the magnitude of the sampling rate. Hence, it is more suitable for data sampled at a low sampling rate. Conversely, the weak formulation is a proficient method for attenuating noise and mitigating noise amplification, but it may also suppress signals at high frequencies. It is suitable for data that is sampled at a rate significantly higher than the Nyquist frequency of the signal. In practice, the sampling rate is often limited by the capabilities of the measuring hardware, and problems that push the limit of the equipment are often sampled at a frequency just above the Nyquist frequency of the relevant dynamics. As a rule of thumb, the weak formulation usually requires at least 6-12 data points in the integral. Hence, practically, it can only resolve dynamics 6-12 times slower than the sampling frequency. By improving model selection performance in the high noise regime, the Bayesian reformulation presented in this study improves the overall chances of recovering the correct model even when the sampling rate is close to the Nyquist frequency of the dynamic, i.e. in the low-data regime.
@@ 3 Examples
This section will showcase the noise robustness and data efficiency of our Bayesian-SINDy algorithm in recovering the desired equations precisely in comparison with the original SINDy algorithm [1] that is based on Sequentially Thresholded Least Squares (STLS). We will also compare our Bayesian-SINDy algorithm with SparseBayes [25] to show the improvements our Bayesian-SINDy has brought to the Bayesian approaches. Four examples will be used to compare the three algorithms, in which we will consider a polynomial library $\Theta(\mathbf{X})$ up to third order. To compare the three algorithms and to ensure the result is not skewed by individual runs, we will vary the sampling frequency and noise variance for the first three synthetic examples, and iterate the learning process 10000 times to compare the percentage of runs for which each algorithm has successfully recovered exactly all the terms in the original equation without extra spurious terms, i.e. the success rate of recovering the original equations. In this study, the use of success rate as the metric is justified because we are focused on the most parsimonious and generalisable equations, which are the original equations, precisely. Furthermore, to avoid miscounting models with the same sparsity pattern but substantially different coefficients, the success rate is only counted when the model coefficient error (MCE = $||\mathbf{w}-\hat{\mathbf{w}}||_{2}/||\hat{\mathbf{w}}||_{2}$, where $\hat{\mathbf{w}}$ is the truth) is below $25\%$. As for the fourth example, we will learn the dynamical equations from real-world measurements. Although a ground-truth model is absent, historical uses of the dataset [38, 39, 32] have shown that we should be expecting the Lotka-Volterra equations.
We should note that in this section, we do not aim to claim that Bayesian-SINDy is the most effective method for recovering the correct dynamical equations from data. Furthermore, we note that SINDy (STLS) is not the most effective methodology among many SINDy extensions. As we shall discuss later in §4.2, we expect more computationally-intensive sampling-based methods to be more robust. However, by comparing SINDy (STLS) with an equivalently simple and computationally efficient method derived from the Bayesian framework, we will demonstrate that a Bayesian approach is more robust in noisy and low data scenarios, and thereby encourage a Bayesian perspective to other extensions of SINDy. The comparison with SparseBayes, meanwhile, demonstrates the improvement we have brought to our learning algorithm. To compare the algorithms properly, we have also taken steps to ensure they are performing the best they can with the appropriate hyperparameters. Details of the hyperparameters used for each algorithm and in each example are listed in Appendix B.
@@ 3.2 Cubic oscillator
Figure 5: The phase diagram of the cubic nonlinear oscillator system (21, black line) and the system estimated by SINDy (STLS) (blue), Bayesian-SINDy (red) and SparseBayes/RVM (yellow) given the noisy data (crosses) of the system. The noise level are $(a)$ $\sigma_{x}=0.005$ and $(b)$ $\sigma_{x}=0.01$ and the data are sampled at $\Delta t=0.05$ for $t\in[0,5]$. The right panel shows the model equations from each algorithm when $\sigma_{x}=0.01$. In $(a)$, the prediction by SINDy and B-SINDy is so close to truth their differences are practically indiscernible.
Figure 6: The percentage of runs that successfully identifies the cubic nonlinear oscillator equations from data using SINDy (STLS, blue), Bayesian-SINDy (red) and SparseBayes (yellow) after 10000 iterations against the number of data points $\tilde{N}=5/\Delta t$. The noise levels are $(a)$ $\sigma_{x}=0.005$ and $(b)$ $\sigma_{x}=0.01$.
@@
Figure 7: Cubic oscillator: breakdown of contributions to the total variance of noise $\boldsymbol{\sigma}$ in the dynamics of $(a)$ $x_{1}$ and $(b)$ $x_{2}$ by each term in (8). Thick lines show the variance computed from the statistics of 1000 iterations, and lighter lines (shaded area) show the bound of the theoretical calculation using (8) from each iteration. Here, $\Delta t=0.1$.
@@ 3.3 Lorenz system
Figure 8: The percentage of runs that successfully identifies the Lorenz system equations from data sampled between $t\in[0,2.5]$ using SINDy (STLS, blue), Bayesian-SINDy (red) and SparseBayes (yellow) after 10000 iterations against the number of data points $\tilde{N}=2.5/\Delta t$. The noise levels are $(a)$ $\sigma_{x}=0.025$, $(b)$ $\sigma_{x}=0.05$, $(c)$ $\sigma_{x}=0.1$ and $(d)$ $\sigma_{x}=0.2$. The dashed line shows the success rate of recovering the model without the $x_{2}$ term in the equation for $\dot{x}_{2}$. The inset in $(d)$ shows the success rate when the time interval is extended to $t\in[0,10]$.
However, in figure 8$(d)$, in which the noise is $\sigma_{x}=0.2$, none of the three algorithms can recover the model exactly. Both SINDy (blue) and SparseBayes (yellow) can only recover the full model sporadically and showed little consistency in the model they produced. In contrast, Bayesian-SINDy consistently recovers a model with only the $x_{2}$ term in equation for $\dot{x}_{2}$ missing (as shown by the dashed red line). The missing $x_{2}$ term in $\dot{x}_{2}$ is particularly difficult to learn because it is highly collinear with the existing $x_{1}$ term in the model. To distinguish the two would require more data in regions in which the two differ significantly, but, as shown in figure 9, the given signal is highly obscured by noise and does not sample those regions sufficiently in the short time interval $t\in[0,2.5]$. In other words, there is simply not enough information in the small time window to infer the $x_{2}$ term in the equation for $\dot{x}_{2}$. From the evidence-maximisation perspective, this difficulty is reflected in figure 1$(b)$, where $\mathcal{J}(\alpha_{m})>0$ only for a small range of $\alpha_{m}$. In practice, when the data is too low or too noisy, information content in the data may not be enough to infer the full dynamics. In such cases, one inevitably has to increase the amount of data available by considering a long time interval, say, $t\in[0,10]$. As demonstrated in the inset of figure 8$(d)$, Bayesian-SINDy can eventually recover the full dynamics most of the time, while SINDy remains less robust in this high noise regime. This lack of information, as depicted here in figure 9, also hints at the possibility of further improving data efficiency if we are allowed to sample data in the most informative state. We will discuss this information-theoretical approach further in §4.3.
Figure 9: Learning the dynamics of $\dot{x}_{2}-rx_{1}+x_{1}x_{3}$ in the Lorenz system: the crosses show the values of $\dot{x}_{2}-rx_{1}+x_{1}x_{3}$ from the data, which should match $x_{2}$ (yellow line). However, $x_{1}$ (red line) is highly collinear with $x_{2}$, and it is not obvious which term should fit the data better. Both $x_{1}$ and $x_{2}$ are scaled to best match the data $\dot{x}_{2}-rx_{1}+x_{1}x_{3}$. Here, $\Delta t=0.025$ ($\tilde{N}=100$) and $\sigma_{x}=0.2$.
@@ 3.4 Lotka-Volterra equation from Lynx-Hare population data
In this example, we apply the three algorithms to model the population dynamics of Lynx and Hare in Canada, using pelt trade data collected by the Hudson Bay Company between 1900 and 1920 [38, 39]. Assuming the pelts of these two species traded by the Company are proportional to the true population, it is expected that the pelt trade data will approximately adhere to the Lotka-Volterra model for predator-prey dynamics. In fact, this dataset has been used as empirical evidence to support the Lotka-Volterra equations, and has become increasingly popular as a real-life benchmarking dataset among statisticians and data scientists [40, 12, 32]. The primary challenge posed by this dataset is its limited number of data points - specifically, 21 annual data points for each population, resulting in a total of 42 data points. Due to the low sampling rate, the finite difference method is more appropriate, but it will, as one expects, also amplify the noise in the derivative, which we will then mitigate by relying on the noise robustness of the Bayesian framework. The weak formulation was also tried, but the low-pass filtering effect filtered vital signals in the data, leading to subpar learning performance across all methods.
@@
To address the difficulty, we ran a hyperparameter sweep on $\sigma_{x}$, assuming $\alpha=10^{-2}$, and chose the value of $\sigma_{x}\approx 2.7$ by evidence maximisation. Although this technique is similar to the evidence maximisation process in SparseBayes, a notable distinction is that our procedure does not assume a uniform $\boldsymbol{\sigma}$ for all data points. Instead, we calculate it using equation (8) and algorithm 1. The parameters estimated are displayed in table 1, where the Lotka-Volterra model is successfully recovered by Bayesian-SINDy, while the other two struggle to learn a consistent model. Uncertainties in the estimated parameters, also shown in table 1, are also easily recovered from the posterior distribution. Figure 10 also shows how the uncertainty in the estimated parameter propagates through the dynamics, assuming the same initial condition. Notably, our method was able to estimate the parameter uncertainties without expensive sampling, unlike previous techniques [32].
@@ 4.1 Scaling of likelihood and Occam factor with $N$ and $\boldsymbol{\beta}$
$\mathcal{J}(\mathbf{y}|\mathbf{D},\mathbf{w},\boldsymbol{\beta})=-\frac{N}{2}\ln{2\pi}-\frac{1}{2}||\boldsymbol{\beta}||_{1}-\frac{1}{2}\boldsymbol{\beta}\cdot(\mathbf{y}-\mathbf{D}\mathbf{w})^{2}.$ (23)
@@
$\mathcal{J}(\mathbf{y}|\mathbf{D},\boldsymbol{\beta},\boldsymbol{\alpha})=\mathcal{J}(\mathbf{y}|\mathbf{D},\mathbf{w},\boldsymbol{\beta})|_{\mathbf{w}=\boldsymbol{\mu}}+\mathcal{J}(\mathbf{w}|\boldsymbol{\alpha})|_{\mathbf{w}=\boldsymbol{\mu}}+\frac{1}{2}\ln{(\det(\boldsymbol{\Sigma}))},$ (24a)
@@
the log-likelihood $\mathcal{J}(\mathbf{y}|\mathbf{D},\mathbf{w},\boldsymbol{\beta})$, and the log-prior $\mathcal{J}(\mathbf{w}|\boldsymbol{\alpha})$ at the maximum a posteriori (MAP) point. The MAP point refers to the parameter values $\mathbf{w}=\boldsymbol{\mu}$ that maximise the posterior.
@@
Figure 11: Difference in log-evidence, log-likelihood and the Occam factor (i.e. log-evidence minus log-likelihood) compared to the highest-ranked model of all possible models for $\dot{x}_{3}$ in the Lorenz system. Models are ranked by $(a,c,e)$ evidence and $(b,d,f)$ likelihood. Insets show a zoomed plot of the first few hundred models, and the barcode plots below show the non-zero terms in each model, marked as green (correct) or orange (incorrect). The correct model is always ranked highest by evidence in $(a,c,e)$, and indicated with a star in $(b,d,f)$. Meanwhile, the model selected by highest likelihood contain all the terms, and the models selected by SINDy are shown in the middle. Here, $(a-b)$ $\sigma_{x}=0.2$ and $\tilde{N}=200$, $(c-d)$ $\sigma_{x}=1$ and $\tilde{N}=200$, and $(e-f)$ $\sigma_{x}=1$ and $\tilde{N}=50$.
It should be noted that in equation (23), the log-likelihood scales linearly with the magnitude of the inverse noise variance $\boldsymbol{\beta}$ and the number of data points $N$. On the other hand, the remaining components in the log-evidence, also known as the Occam factor, serve as a penalty for models that have an excessive number of parameters. The Occam factor scales with the number of terms $M$ in the model and only has a partial and weak $\ln{N}$ scaling with $N$.
@@
Figure 11 demonstrates how the scaling of the log-likelihood and the Occam factor plays out in the model selection process. The left column $(a,c,e)$ of figure 11 shows the contribution of the log-likelihood (blue) and the Occam factor (red) to the log-evidence across all 1024 possible models for $\dot{x}_{3}$ in the Lorenz system when a smaller, second-order polynomial library containing ten terms is considered. For each model, the barcode plot below show the library terms that are non-zero, with the non-zero terms $x_{3}$ and $x_{1}x_{2}$ in the true model coloured green. The models are sorted by log-evidence, demonstrating a Bayesian perspective on model selection, in which the model with maximum evidence is selected as the most parsimonious model for the data. The maximum evidence model (Bayesian-SINDy) identifies the correct model in all three scenarios.
For the same dataset, the right column $(b,d,f)$ of figure 11 sorts the model based on log-likelihood calculated from the maximum likelihood estimate (MLE), i.e. $\mathbf{w}=\boldsymbol{\mu}_{\text{MLE}}$, which can be obtained by ordinary least squares. (Note that when $N$ and $\beta$ is large, the MLE and MAP estimates of $\boldsymbol{\mu}$ converge.) In contrast to the Bayesian perspective on the left, the right column reflects a frequentist perspective in which the model with the best fit is selected as the best model for the data. As one would expect, the MLE-based model selection always selects the model with all ten terms (i.e. over-fit). To induce sparsity, the frequentist approach STLS introduces an empirical thresholding process. The idea is to remove small coefficient terms that don’t contribute much to the likelihood, thereby achieving a compromise between maximising likelihood and sparisty.
@@
When the data is abundant and the noise is minimal (i.e. $N||\boldsymbol{\beta}||\gg M$), models that include the correct terms have significantly higher likelihood and are easily distinguishable from the other models when sorted by likelihood. This is demonstrated by the continuous green band in the barcode plot in 11$(b)$. Phenomenologically, this signifies that all models containing the appropriate terms exhibit comparable levels of likelihood. STLS leverages this clear separation in likelihood and only eliminates irrelevant terms that do not affect the likelihood much in its thresholding process, thereby effectively constraining its search for the most sparse model within this region of large likelihood.
@@
However, as the amount of data is lowered (i.e. smaller $N$), or when the noise increases (i.e. smaller $||\boldsymbol{\beta}||$), the distinct separation of the high likelihood region from the other possible models becomes less evident. In these scenarios, models lacking the relevant terms are more likely to fit the data incidentally, as reflected by the discontinuous green band in the barcode plot of figure 11$(d,f)$. STLS may, thus, converge on an incorrect model. Alternatively, STLS may not converge to the most sparse model because there is a non-negligible chance that irrelevant terms will also have coefficients larger than the threshold when $N$ or $\beta$ decreases.